OpenStudy (anonymous):

Are the variables in a direct or inverse variation equation interchangeable? Like, if P1/v1=p2/v2, does that mean that v1/p1=v2/p2 as well?

OpenStudy (anonymous):

i'm not going to ask for answers or anything like that, I just need to know the theory.

OpenStudy (dtan5457):

I believe so. Here's an example (gay-lussac's law) p1/t1=p2/t2 "The pressure exerted by an ideally behaving gas is 7 kilo pascals at 200k. What pressure is the gas at 500k? normally 7/200k=p2/500k same thing as 200k/7=500k/p2 both=17.5 kpa

OpenStudy (anonymous):

It is very correct @coolshava . That is because they still vary directly proportional. Just take this example: